OpenStudy (anonymous):

A ladder is resting against a wall. The ladder and the ground make an angle of 45° and the ladder is 7 ft from the wall. How long is the ladder?

OpenStudy (anonymous):

Just move all values to one side and leave the other side to zero and there's your quadratic formula. Let x be the distance from the wall to the bottom of the ladder Let 2x+1 be the length of the ladder 15^2 + x^2 = (2x+1)^2 4x^2 + 4x + 1 = 15^2 + x^2 I just expanded and switched the sides 3x^2 + 4x - 224 = 0 x^2 + 4x - 672 = 0 <-- bridge method is not in the NYC curriculum but I don't know about yours (x+28)(x-24) = 0 (3x+28)(3x-24) = 0 (3x+28)(x-8) = 0 x = 8

OpenStudy (anonymous):

8.76 FT?